Country Z has an economy that can be described as

C= 200 + (0.8) y

I= 150

G= 250

X= 115

M= 125 + (0.05) Y

C= Consumption

Y= Income

I= investment

G= Goverment spending

X= exports

M=Imports

What is the equilibrium level of income for this economy?

1. 2360

2. 3930

3. 4,960

4. 8,260
Answer #1

Answer

The equilibrium is at

Y=C+I+G+X-M

Y=200+0.8Y+150+250+115-(125+0.05Y)

Y=590+0.8Y-0.05Y

Y-0.75Y=590

0.25Y=590

Y=2360

The equilibrium is at $2360

Option 1